Brenda Chatterjee is wealthy, single, and generous. During the previous year she made the following gifts:

•Gift of stock to the local art museum (a 501(c)(3) nonprofit organization): $245,000

•College tuition payment for niece: $29,000 made directly to the institution

•Medical bills for elderly neighbor: $18,000 paid directly to the hospital

•Home down payment gift to daughter: $30,000 paid directly to the mortgage lender

•Cash gift to son: $17,000Plus, Youll Pay This Percentage on the Amount in Excess of the Lower Limit For Taxable Estates Between ... Youll Pay This A

•Use the gift tax rates shown below as a guide to answer the following questions:

a.What is the total amount of taxable gifts in 2018?

b.What is Brenda’s gift tax liability in 2018?

c.Assuming a unified credit of $4,425,800, what alternative does Brenda have in terms of paying the gift tax liability?
Answer #1

a) From the above table we can see that upto $ 9,999, gifts are not taxable. In the question given, all individual gifts are above $ 9,999. Hence, all gifts are taxable.

Therefore, the total amount of taxable gifts in 2018 is:

Gift of stock to local Art museum $ 2,45,000
College Tuition payment for niece $ 29,000
Medical bills for elderly neighbour $ 18,000
Home down payment gift to daughter $ 30,000
Cash gift to son $ 17,000
TOTAL AMOUNT OF TAXABLE GIFTS $ 3,39,000

b) Calculation of Brenda's Tax liability in 2018:

All amount in ($)

PARTICULARS GIFT AMOUNT (A)

LIMIT USED FROM THE ABOVE TABLE (B)

AMOUNT IN EXCESS OF LOWER LIMIT (C) =

(A)- LOWER LIMIT

TAX AMOUNT (D)

EXTRA TAX AS PERCENTAGE OF AMOUNT IN EXCESS OF LOWER LIMIT (E) =

(C)* TAX RATE GIVEN

TOTAL TAX LIABILITY= (D+E)

Gifts to local Musuem 2,45,000 1,50,000 -2,49,999 95.000 38,800

=(95,000*0.32)

=30,400

69,200
College tuition payment for niece 29,000 20,000- 39,999 9,000 3,800

= (9,000*0.22)

=1,980

5,780
Medical bills for elderly neighbour 18,000 10,000- 19,999 8,000 1,800

= (8,000*0.20)

=1,600

3,400
Home down payment gift to daughter 30,000 20,000- 39,999 10,000 3,800

= (10,000*0.22)

=2,200

6,000
Cash gift to son 17,000 10,000- 19,999 7,000 1,800

= (7,000*0.20)

=1,400

3,200
87,580

Hence, the total tax liability of Brenda in 2018 is $ 87,580.

c). Unified Credit is the total amount upto which an individual is allowed to give gifts to any person without paying any tax. Once the aggregate amount of gifts breach that limit, an individual will be liable to pay tax on the amount of gifts given above that limit.

In the above question, assuming that Brenda is having a unified credit of $ 4,425,800 and till now she has not breached this limit. Therefore, She will not be liable to pay any gift tax as the total amount of taxable gifts is $ 3,39,000.

Until and unless, the aggregate value of gifts given by her in lifetime does not breach the limit of $ 4,425,800, Brenda will not be liable to pay any tax on such gifts.